On January 1, Year 1, Hills purchased equipment for $350,000. The equipment had an estimated useful life of ten years and an estimated residual value of $50,000. Use the double-declining-balance method to answer the following question(s).

Refer to Exhibit 11-1. How much depreciation should Hills record on the asset in year 2?

16,000

24,000

56,000

48,000
Answer #1

Double declining rate = (2/useful life) = (2/10) = 20%

Depreciation for year 1 = 350,000 * 20% = 70,000

Depreciation for year 2 = (350,000-70,000)*20% = 56,000

Option C is the answer